Two small planes start from the same point and fly in opposite directions. The first plane is flying 40mph slower than the second plane. In the 3 hr the planes are 1920 mi apart. Find the rate of the faster plane.
Answer by richwmiller(17219)

3*(x-40)+3x=1920
3x-120+3x=1920
6x=2040
x=340
x-40=300